OpenStudy (anonymous):

Which equation could have been used to create this function table? A. y = 4x B. y = 5x C. y = x + 7 D. y = x + 2

OpenStudy (flower2468):

what do you thiink?

OpenStudy (anonymous):

I don't even no how to do this

OpenStudy (flower2468):

okay so y= the numbers on the right the y colum and the equaions mean like so choice a is =y=4x, so take the first number on the x side and plug it in the equation= y=4*6

OpenStudy (anonymous):

so 4*6=24

OpenStudy (flower2468):

wait no, 4*4=

OpenStudy (flower2468):

not 6, I i meant 4 oops

OpenStudy (anonymous):

so it would be 16

OpenStudy (flower2468):

and is that the number on the other side on 4 in the table?

OpenStudy (anonymous):

6

OpenStudy (flower2468):

no the number 16 is not which means that a is wrong

OpenStudy (anonymous):

okay

OpenStudy (flower2468):

now lets try B now plug in 4 in the eqation: y=5*4

OpenStudy (anonymous):

So it must be b

OpenStudy (flower2468):

well is 20, which is 5*4 on the other side in the y colum?

OpenStudy (anonymous):

idk

OpenStudy (flower2468):

it is not, so I already tell you it is not c, because 4+7 does not=6

OpenStudy (flower2468):

so for D, 4+2=6 so that is right and 5+2=7 and 7+2=9 and 10+2=12 and 11+2=13

OpenStudy (flower2468):

so d is right because the rule is add2

OpenStudy (flower2468):

see how i did that?

OpenStudy (anonymous):

yes